# Uncertainty Quantification in Retrieval Augmented Question Answering

### v2

@@ 6 Conclusions
In this work we address uncertainty estimation in the context of retrieval augmented QA with a method that relies on individual passage utilities. Key in our approach is the definition of utility in terms of the behaviour of the QA model and whether it is able to provide a correct answer given a retrieved passage. We train a small neural model on passage utility judgements elicted from the QA model’s responses and use utility predictions to estimate answer uncertainty. Experimental results show that our uncertainty estimator is competitive or better than existing strong methods while being light-weight. Future work could extend this approach to long-form generation tasks (Stelmakh et al., 2022; Gao et al., 2023; Min et al., 2023; Zhang et al., 2024a) where evaluating whether answer correctness is more challenging (Zhang et al., 2024b) and to multi-modal QA scenarios (Borszukovszki et al., 2025).
Appendix A Ablation of Passage Utility Predictor Training Objective
@@
Table 3 shows AUROC results on answer uncertainty prediction with Passage Utility estimators trained with different variants of the objective in Equation (5). The first row shows the full objective (see training details in Section E.2), the second row shows a variant where the ranking objective uses only entailment annotations ($e$), and in the third row the objective is solely based on accuracy prediction ($\mathcal{L}_{BCE}$). As can be seen, there is a drop in performance when the pairwise ranking loss is not used (i.e., last line of Table 3); this component of the objective provides a smoother signal on passage utility which is empirically beneficial. However, when the pairwise ranking loss is only based on entailment, performance drops by several points, highlighting the importance of combining both to model the QA answering behaviour.
@@ Appendix C Generality of the Uncertainty Aggregation Strategy
To validate the intuition behind our uncertainty estimation step (Section 3.2), we compare the behaviour of the QA model when prompted with individual passages in $R$ versus when prompted with $R$ (i.e., the top-$|R|$ passages). In particular, we want to inspect the proportion of cases where our stratege of taking the maximum uitility score among the passages in $R$ does not agree with the entire set accuracy. In other words, we are interested in cases where the QA model is accurate with at least one individual passage (IndAcc\,$\uparrow$) but answers incorrectly when promtped with $R$ and vice versa. In this study, we consider two individual passage accuracy variants (related to the combined definition of Passage Utility Section 3.1). One is based on accuracy $a$ (see Section 4) and is either 0 or 1. The other one is also based on accuracy but smoothed by entailment $e$ (computed by an off-the-shelf entailment model; see Section 4) and downgrades cases where $a=1$ into $a=0$ if $e<0.5$. The latter occurs in cases where the QA model produces an answer that is accurate but not entailed by the passage.
We analyse the behaviour of four QA models (LLMs) across five datasets in our experimental setup. Table 5 shows the proportion of model disagreements across development sets. We can see that such disagreements amount to a relatively small number in both settings, i.e., when at least one individual passage in $R$ yields a correct answer (IndAcc\,$\uparrow$) but the QA model prompted with $R$ yields an incorrect answer (RAcc\,$\downarrow$) and vice versa (RAcc\,$\uparrow$ IndAcc\,$\downarrow$). The results in Table 5 confirm that our aggregation approach based on individual passages is fairly general. It approximates answer uncertainty when prompting with $|R|$ passages while avoiding the complexity of estimating uncertainty over all possible combinations of input passages in terms of number and order.
@@
RAcc\,$\uparrow$ IndAcc\,$\downarrow$ $a$ 0.01 0.01 0.01 0.01
Table 5: Average proportion of cases (five development sets) where at least one individual passage in $R$ leads a correct answer (IndAcc\,$\uparrow$) but the QA model prompted with $R$ yields an incorrect one (RAcc\,$\downarrow$) (and vice versa RAcc\,$\uparrow$ IndAcc\,$\downarrow$).
@@ Appendix D Out-of-distribution Generalization of Uncertainty Estimation
We next assess the generalization ability of our Passage Utility estimator. Following previous work on question answering and out-of-distribution (o.o.d) scenarios (Kamath et al., 2020; Zhang et al., 2021), we train a Passage Utility predictor on the SQuAD dataset and then use it to predict zero-shot (i.e., without further fine-tuning) passage utilities on all other datasets (test set). As p(true) relies on 20 in context training examples, we also evaluate its ability to generalise in out-of-distribution settings.
@@
Table 6 shows AUROC for answer uncertainty prediction in o.o.d scenarios. As an upper bound, the upper-block of the table shows AUROC values for the in distribution (i.i.d) scenario for Passage Utility and p(true). We compare o.o.d performance w.r.t. PPL and Semantic Entropy which do not rely on training examples. Although Passage Utility performance decreases in o.o.d settings, it remains competitive in four out of five datasets. In these four cases, it is always statistically significantly different from the PPL method and comparable to p(true) and Semantic Entropy. Interestingly, p(true)’s performance also drops in all o.o.d test sets showing that relying on a fixed number of in-context learning examples is neither a robust nor scalable adaptation method. Note that we focus on zero-shot accuracy to assess bare transfer performance. However, it would make sense to train a unified Passage Utility predictor for all QA tasks following a meta-learning scheme for better adaptation (Finn et al., 2017). In addition, the model could be continually adapted via few examples (Kamath et al., 2020; Zhang et al., 2021). However, we leave this to future work.
NQ TQA WebQ PopQA RefuNQ
p(true) (i.i.d) 0.73 0.75 0.67 0.81
Passage Utility (i.i.d) 0.76 0.85 0.69 0.86
PPL 0.64   0.68 0.52   0.59 0.51
@@
p(true) (o.o.d) 0.67 0.63 0.63 0.72 0.62
Passage Utility (o.o.d) 0.65 0.79 0.60 0.72 0.79
Table 6: Out-of-domain performance of Passage Utility predictor for Gemma2-9B (lower block). Best values are highlighted in bold; we also mark with  next best values which are significantly different using the paired De Long test ($p<0.05$). Uncertainty predictors are trained on SQuAD and evaluated zero-shot on NQ, TQA, WebQ, PopQA, and RefuNQ test sets. The upper block reports AUROC values from our main in-distribution experiments (Table 1).
Appendix E Experimental Details
E.1 Datasets and Splits
@@ E.2 Implementation Details
Passage Utility Predictor
@@
In this section, we describe additional answer uncertainty estimation methods (for which we present supplementary results in Section F). Maximum Sequence Probability (MSP) is based on the probability of the most likely answer and is computed as:
@@
Note that, in contrast to $\text{PPL}(x,R,\mathcal{M})$ reported in the main section of the paper, this metric is biased by answer length, i.e., identifying an answer to have low probability (low confidence) because of its length. Despite the fact that QA models are instructed to produce short answers, they do not always follow instructions. For this reason, we consider perplexity a more accurate metric. Indeed, answer length could indicate that the model is uncertain about the answer. Thus, we also estimate answer uncertainty from the Average Answer Length (AvgAnsLen) as the average number of words in the sampled answers. As seen in Section F.2, Table 14, MSP and AvgAnsLen perform similarly across the board.
@@
We use the implementation provided in Farquhar et al. (2024) to compute AUROC, Accuracy at Xof rejection, and AURAC metrics.
@@
p(true) prompt Question: question Brainstormed Answers: most likely answer sampled answer 1 sampled answer N Possible answer: most likely answer Is the possible answer: A) True B) False The possible answer is: correct choice Knowledge: [1] passage [2] passage [$|R|$] passage Question: question Brainstormed Answers: most likely answer sampled answer 1 sampled answer $N$ Possible answer: most likely answer Is the possible answer: A) True B) False The possible answer is:
Table 10: Prompt used for p(true) approach. Items in blue are filled with in-context examples from the training set and the current example being evaluated. $N$ represents the number of sampled answers. The “sequence of in-context examples” prefix is a sequence of examples taken from the training split with the same question format but with the answer to The possible answer is: resolved.
Accuracy evaluation prompt. You need to check whether the prediction of a question-answering system to a question is correct. You should make the judgment based on a list of ground truth answers provided to you. Your response should be "correct" if the prediction is correct or "incorrect" if the prediction is wrong. Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: Roma Gill and W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"]" Prediction: Roma Shakespeare Correctness: incorrect Question: What country is Maharashtra Metro Rail Corporation Limited located in? Ground truth: ["India"] Prediction: Maharashtra Correctness: incorrect Question: What’s the job of Song Kang-ho in Parasite (2019)? Ground truth: ["actor"] Prediction: He plays the role of Kim Ki-taek, the patriarch of the Kim family. Correctness: correct Question: Which era did Michael Oakeshott belong to? Ground truth: ["20th-century philosophy"] Prediction: 20th century." Correctness: correct Question: Edward Tise (known for Full Metal Jacket (1987)) is in what department? Ground truth: ["sound department"] Prediction: 2nd Infantry Division, United States Army Correctness: incorrect Question: What wine region is Finger Lakes AVA a part of? Ground truth: ["New York wine"] Prediction: Finger Lakes AVA Correctness: incorrect Question: question Ground truth: gold answer Prediction: generated answer Correctness:

### v3

@@ 7 Conclusions
In this work we address uncertainty estimation in the context of retrieval augmented QA with a method that relies on individual passage utilities. Key in our approach is the definition of utility in terms of the behaviour of the QA model and whether it is able to provide a correct answer given a retrieved passage. We train a small neural model on passage utility judgements elicited from the QA model’s responses and use utility predictions to estimate answer uncertainty. Experimental results show that our uncertainty estimator is competitive or better than existing strong methods while being light-weight. Future work could extend this approach to long-form generation tasks (Stelmakh et al., 2022; Gao et al., 2023; Min et al., 2023; Zhang et al., 2024a) where evaluating whether answer correctness is more challenging (Zhang et al., 2024b) and to multi-modal QA scenarios (Borszukovszki et al., 2025).
Acknowledgments
We thank the action editor and anonymous reviewers for their constructive feedback. We want to thank Ivan Titov and Gustavo Giménez for useful discussions; and Alex Yu for sharing the synthetic multi-hop QA scripts. We are also grateful to Harry Sfyrakis and the UK Research and Innovation via the Innovate UK award (10092998) to Algomo Limited. This work was supported by the UK Engineering and Physical Sciences Research Council (grant EP/W002876/1). We also gratefully acknowledge the support of the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.
Appendix A Ablation of the Passage Utility Predictor Training Objective
@@
Table 4 shows AUROC results on answer uncertainty prediction with Passage Utility estimators trained with different variants of the objective in Equation (5). The first row shows the full objective (see training details in Section F.2), the second row shows a variant where the ranking objective uses only entailment utility annotations ($e$), and in the third row the objective is solely based on accuracy prediction ($\mathcal{L}_{BCE}$). As can be seen, there is a drop in performance when the pairwise ranking loss is not used (i.e., last line of Table 4); this component of the objective provides a smoother signal on passage utility, which is empirically beneficial. However, when the pairwise ranking loss is only based on entailment, performance drops by several points, highlighting the importance of combining both to model QA answering behaviour.
Table 5 reports various ablation studies (with Gemma2-9B on the WebQ development set) with different instantiations of the Passage Utility score $\upsilon_{\mathcal{M}}$ and our training objective. We present different ways to combine accuracy ($a$) and entailment ($e$) scores to induce utility rankings with the pairwise ranking loss $\mathcal{L}_{rank},(\cdot)$ (Equation 3). These include average $(a+e)/2$ as in Equation 2, addition $(a+e)$, addition but inverting the entailment score ((e|1e)+a)((e|1-e)+a) when the passage yields an inaccurate answer ($a=0$), or when one of the two is given a zero weight, i.e., $(a)$ or $(e)$ alone. We asses these ranking variants when using the ranking objective alone (rightmost block) as well as when combined with the binary cross-entropy objective $\mathcal{L}_{BCE}$ (Equation 5) (middle block); we also report performance when only the $\mathcal{L}_{BCE}$ objective is used (leftmost block). All model variants are trained with rank-based model selection and $\lambda=0.25$ following training details in Section F.2.
When looking at the pairwise ranking objective alone (rightmost block), entailment dominates the ranking and the utility scores learnt; $(e)$ pairwise ranking as well as $(a+e)$ variants yield utility scores that have similar discriminative power. The $(a)$ variant exhibits the worst performance, given that many pairs are discarded due to having the same utility. AUROC values improve when the pairwise ranking is combined with the cross-entropy objective (middle block). In this case, in addition to enforcing the pairwise ranking with $\mathcal{L}_{rank},(\cdot)$, the utility scores are regularised. In other words, the utility scores of a pair of passages with the same accuracy (i.e., only ordered by entailment) will end up closer (differ less) than the utility scores of a pair of passages with different accuracies (differ more). This is reminiscent of the Bradley-Terry model (Bradley & Terry, 1952): in the first case, one passage is better than the other with a small probability, while in the second case the probability of one being better than the other is higher. The exception here is the $\mathcal{L}_{rank},(e)+\lambda\,\mathcal{L}_{BCE}$ variant where the pairwise ranking by entailment may often contradict the binary cross-entropy signal.
@@
$\mathcal{L}_{BCE}$ $\mathcal{L}_{rank}+\lambda\,\mathcal{L}_{BCE}$ $\mathcal{L}_{rank}$
$(e+a)/2$ $(e+a)$ ((e|1e)+a)((e|1-e)+a) $(e)$ $(e+a)/2$ $(e+a)$ ((e|1e)+a)((e|1-e)+a) $(e)$ $(a)$
0.71 0.75 0.74 0.72 0.63 0.65 0.63 0.62 0.61 0.53
Table 5: Answer uncertainty estimation with Passage Utility predictors trained with different variants of the objective in Equation (2). AUROC values are for Gemma2-9B and the WebQ development set.
@@ Appendix C Generality of the Uncertainty Aggregation Strategy
[…]
@@
[…]
@@ Appendix D Out-of-distribution Generalization of Uncertainty Estimation
We assess the generalization ability of our Passage Utility estimator both in terms of new QA tasks and QA models. We train a unique Passage Utility predictor for the Gemma2-9B model. Following previous work on question answering and out-of-distribution (o.o.d) scenarios (Kamath et al., 2020; Zhang et al., 2021), we train it on the SQuAD dataset and then use it to predict zero-shot (i.e., without further fine-tuning) passage utilities on all other datasets (test set). As p(true) relies on 20 in context training examples, we also evaluate its ability to generalise in out-of-distribution settings.
@@
Table 8  (QA Task block) shows AUROC for answer uncertainty estimation in o.o.d scenarios. As an upper bound, the i.i.d block of the table shows AUROC values in the in-distribution scenario for Passage Utility and p(true). We compare o.o.d performance w.r.t. PPL and Semantic Entropy which do not rely on training examples. Although Passage Utility performance decreases in o.o.d settings, it remains competitive in four out of five datasets. In these four cases, it is always statistically significantly different from the PPL method and comparable to p(true) and Semantic Entropy. Interestingly, p(true)’s performance also drops in all o.o.d test sets showing that relying on a fixed number of in-context learning examples is neither a robust nor scalable adaptation method.
To understand the observed performance drop, we conducted a comparative analysis of passage utility predictions for 50 question-passage pairs. We examined predictions from a utility predictor trained on WebQ data (i.i.d) versus one trained on SQuAD (o.o.d). We sampled 10 WebQ test questions: 5 for which the predicted utility decreased from high under i.i.d. conditions to low under o.o.d. conditions (changing from true negative to false positive), and 5 for which it increased in the opposite direction (changing from true positive to false negative). In both scenarios, the o.o.d predictions were predominantly influenced by token overlap and similarity. We hypothesize that when faced with o.o.d questions (e.g., in terms of type, length, or topic), the prediction mechanism defaults to predictions based on a general notion of question-passage similarity, disregarding whether the QA model can answer the question with the given passages. For passages with low (high) similarity to the question, it will predict low (high) utility scores. This behaviour may be a consequence of under-training, suggesting that the Passage Utility predictor cannot predict the accuracy of the target QA model on these o.o.d questions and their corresponding retrieved passages.
To empirically validate these observations, we calculate Spearman’s $\rho$ correlation coefficients between the predicted Passage Utility scores and the corresponding retriever relevance scores (as computed by Contriever-MSMARCO; Izacard et al. 2022) for each passage, examining correlations separately for both i.i.d and o.o.d predictions. We use the Retriever Score because it measures the semantic relation between questions and retrieved passages. In agreement with our manual inspection, there is a positive correlation between Passage Utility and Retriever Score in the o.o.d setting (0.536 with $p$-value < 0.01), suggesting that Passage Utility falls back to textual similarity. In contrast, the i.i.d setting exhibits a negative and weaker correlation (-0.318 with $p$-value < 0.05), indicating that training successfully aligns Passage Utility predictions with the target QA model’s actual performance rather than superficial retriever similarity scores.
We also evaluate generalisation to a new QA model, by training the Passage Utility predictor on utility labels observed for Gemma2-9B and use its predictions to estimate uncertainty for the bigger Gemma2-27B model. We also evaluate p(true) in this o.o.d setting, i.e., we create p(true) prompts as usual with Gemma2-27B generated answers (greedy and sampling), but ask Gemma2-9B to judge the probability of the most likely answer being true. Table 8  (QA Model block) shows AUROC results; for reference, we include in-distribution values (i.i.d block) and compare the o.o.d results with comparison methods. Passage Utility outperforms all other methods across the board, and while still competitive, p(true) exhibits a higher decrease in performance. These preliminary results suggest that in the context of retrieval augmented QA, models behave alike (also suggested by the distribution of observed correct/incorrect individual passage utilities in Table 12). This highlights practical benefits of our approach, such as training a base Passage Utility predictor using data generated by a less expensive model or developing a more general predictor applicable across multiple QA models.
Gemma2-9B NQ TQA WebQ PopQA RefuNQ
QA Task i.i.d p(true) 0.73 0.75 0.67 0.81
Passage Utility 0.76 0.85 0.69 0.86
o.o.d PPL 0.64  0.68 0.52  0.59 0.51
@@
p(true) (SQuAD) 0.67 0.63 0.63 0.72 0.62
Passage Utility (SQuAD) 0.65 0.79 0.60 0.72 0.79
Gemma2-27B NQ TQA WebQ PopQA RefuNQ
QA Model i.i.d p(true) 0.77 0.83 0.67 0.79 0.60
Passage Utility 0.73 0.82 0.69 0.87 0.80
o.o.d PPL  0.64 0.50 0.53 0.53 0.51
Semantic Entropy 0.68 0.62 0.59 0.66 0.58
p(true) (Gemma2-9B) 0.73  0.71  0.64  0.75  0.59
Passage Utility (Gemma2-9B) 0.75 0.80 0.68 0.87 0.79
Table 8: Out-of-domain performance of Passage Utility predictor for Gemma2-9B both in terms of QA task and QA model. i.i.d blocks report AUROC values from our main in-distribution experiments (Table 1); o.o.d blocks contain the o.o.d comparison. Best values are highlighted in bold; we also mark with  next best values which are significantly different using the paired De Long test ($p<0.05$). For the QA task, Passage Utility and p(true) are supervised with SQuAD data and evaluated on NQ, TQA, WebQ, PopQA, and RefuNQ test data. For the QA model, Passage Utility predictors are trained on Gemma2-9B and used to estimate uncertainty for Gemma2-27B; p(true) assessment is provided by Gemma2-9B.
Finally, we train a unique Passage Utility predictor for all QA tasks and assess its generalisation capabilities. To this end, we train a predictor for the Gemma2-9B model on a random sample drawn from the five training sets of size equivalent to the training sets of the individual QA task predictors. Specifically, we took 10k from NQ, TQA, and SQuAD, 3k from PopQA, all WebQ, totalling 35,474 instances for training and 500 instances from each for validation; the number of pairwise instances in the final curated training set is 354,379. For p(true), we mixed 4 randomly sampled examples from each dataset as in-context training examples. We follow the same training procedure as described in Section F.2 (with combined model selection and $\lambda=1$). We show AUROC results in Table 9. Across the six test sets, the unique Passage Utility predictor trained on a mix of QA tasks (bottom block) achieves similar performance to the individual predictors trained on per-task datasets (top block). The unique predictor keeps comparable or better performance than p(true) and outperforms PPL and Semantic Entropy. This preliminary study suggests that is feasible to train a more general predictor for various QA tasks.
NQ TQA WebQ SQuAD PopQA RefuNQ
Task p(true) 0.73 0.75 0.67 0.68 0.81 0.62
Passage Utility 0.76 0.85 0.69 0.80 0.86 0.79
Unique PPL  0.64 0.68 0.52 0.59 0.59 0.51
Semantic Entropy 0.70  0.73  0.58  0.63 0.73 0.59
p(true) 0.71 0.70 0.68  0.63  0.75  0.66
Passage Utility 0.69 0.83 0.68 0.79 0.85 0.79
Table 9: Generalisation of a Unique Passage Utility predictor for Gemma2-9B trained on a mix of QA tasks and evaluated on the six test sets (lower block). Best values are highlighted in bold; we also mark with  next best values which are significantly different using the paired De Long test ($p<0.05$). The upper block reports AUROC values for the Passage Utility from predictors trained for Gemma2-9B on indiviual QA tasks’ training datasets. We also compare p(true) when in-context learning examples are all from the same QA task (upper block) versus from a mix of tasks (lower block).
Appendix E Synthetic Qualitative Analysis of Passage Utility in Multi-hop QA
In multi-hop QA Passage Utility may fail at cases where the QA model cannot answer when prompted with any individual ’hop’ passage, but answers correctly when prompted with the full set. To analyse this limitation, we carry out a qualitative analysis. As discussed in Section 5, there are two scenarios in retrieval augmented multi-hop QA (given existing datasets and current LLMs). Either models can answer with one passage (thus no multi-hop) or retrieval completely fails. Thus, to be able to pinpoint actual multi-hop QA cases and expose the limitations of our approach, we make use of synthetic HotPotQA data (Yang et al., 2018).
First, we use the sets of gold passages provided in the HotPotQA dataset to categorise question types as follows. We prompt the QA model with individual gold passages and with combinations thereof, measuring accuracy. Out of the 500 test questions, 475 are answerable with only one gold ’hop’ passage, 13 require multiple passages, and 12 are unanswerable. Secondly, we focus on the 13 requiring multi-hop questions and insert gold ’hop’ passages within the retrieved set for those where retrieval completely fails. Then, we run QA and uncertainty estimation. In 6 of 13 cases, our approach fails to predict that the QA model’s answer will be correct. It is important to note, however, that these 13 cases represent only 2.6% of the total (13 out of 500).
Appendix F Experimental Details
F.1 Datasets and Splits
@@ F.2 Implementation Details
Curated Passage Utility Dataset
We train our passage utility predictor on a dataset $D_{\mathcal{M}}$ curated from benchmark $D$, e.g., WebQ, consisting of question and gold answer pairs $(x,y)$. For each question we retrieve the top-$k$ passages. Then, we pair question $x$ and retrieved passages $p$ with utility scores $\upsilon_{\mathcal{M}}$ which we obtain after running the QA model  $\mathcal{M}$ on inputs (x,p){(x,p)} and computing the generated answer accuracy and entailment scores (Section 3.1), i.e., we create tuples $(x,p,\upsilon_{\mathcal{M}})$}. From the set of $k$ tuples for question $x$, we derive $k\choose 2$} instances for our pairwise ranking loss.
In experiments, we use $k=5$ retrieved passages per question. Table 11 reports the size (number of training instances) of the curated datasets for each QA task and model. From each question and set of top-$5$ retrieved passages, we derive 10 pairwise ranking instances, discarding those that have equal utilities (e.g., from the WebQ training split with 2,474 question-answer pairs, we curate 24,720 instances with Mistral-7B-v0.3). As our top-$5$ passages are obtained via a real retrieval module, i.e., not synthetically assembled, there are questions for which all passages in the top-$5$ set lead to a correct (incorrect) answer. In these cases, the pairwise ranking is dominated by the entailment score (i.e., accuracy is the same). Table 12 shows the distribution of questions with all retrieved passages leading to the same accuracy (Correct/Incorrect) or mixed (Mixed) accuracies in the curated dataset for each QA task and model.
Models NQ TQA WebQ SQuAD PopQA
Gemma2-9B 395,438 393,475 24,721 393,285 99,770
Gemma2-27B 395,426 393,477 24,723 393,293 99,778
Llama-3.1-8B 790,862 393,465 24,713 393,288 99,787
Mistral-7B-v0.3 395,397 393,474 24,720 393,271 99,772
Table 11: Number of pairwise training instances in the curated datasets to train the Passage Utility predictor with the combined pairwise ranking and binary cross-entropy losses (Section 5).
Incorrect Mixed Correct
Gemma2-9B 21% 52% 27%
Gemma2-27B 20% 53% 27%
Llama-3.1-8B 22% 55% 23%
Mistral-7B-v0.3 21% 56% 23%
Table 12: Number of training instances in curated datasets to train the passage utility predictor with the combined pairwise ranking and binary cross-entropy losses (Section 5).
Passage Utility Predictor Training Details
@@
In this section, we describe additional answer uncertainty estimation methods (for which we present supplementary results in Section G). Maximum Sequence Probability (MSP) is based on the probability of the most likely answer and is computed as:
@@
Note that, in contrast to $\text{PPL}(x,R,\mathcal{M})$ reported in the main section of the paper, this metric is biased by answer length, i.e., identifying an answer to have low probability (low confidence) because of its length. Despite the fact that QA models are instructed to produce short answers, they do not always follow instructions. For this reason, we consider perplexity a more accurate metric. Indeed, answer length could indicate that the model is uncertain about the answer. Thus, we also estimate answer uncertainty from the Average Answer Length (AvgAnsLen) as the average number of words in the sampled answers. As seen in Section G.3, Table 20, MSP and AvgAnsLen perform similarly across the board.
@@
We use the implementation provided in Farquhar et al. (2024) to compute AUROC, Accuracy at X% of rejection, and AURAC metrics.
@@
p(true) prompt Question: question Brainstormed Answers: most likely answer sampled answer 1 sampled answer N Possible answer: most likely answer Is the possible answer: A) True B) False The possible answer is: correct choice Knowledge: [1] passage [2] passage [$|R|$] passage Question: question Brainstormed Answers: most likely answer sampled answer 1 sampled answer NN Possible answer: most likely answer Is the possible answer: A) True B) False The possible answer is:
Table 15: Prompt used for p(true) approach. Items in blue are filled with in-context examples from the training set and the current example being evaluated. NN represents the number of sampled answers. The “sequence of in-context examples” prefix is a sequence of examples taken from the training split with the same question format but with the answer to The possible answer is: resolved.
Accuracy evaluation prompt. You need to check whether the prediction of a question-answering system to a question is correct. You should make the judgment based on a list of ground truth answers provided to you. Your response should be "correct" if the prediction is correct or "incorrect" if the prediction is wrong. Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"] Prediction: Roma Gill and W Shakespeare Correctness: correct Question: Who authored The Taming of the Shrew (published in 2002)? Ground truth: ["William Shakespeare", "Roma Gill"]" Prediction: Roma Shakespeare Correctness: incorrect Question: What country is Maharashtra Metro Rail Corporation Limited located in? Ground truth: ["India"] Prediction: Maharashtra Correctness: incorrect Question: What’s the job of Song Kang-ho in Parasite (2019)? Ground truth: ["actor"] Prediction: He plays the role of Kim Ki-taek, the patriarch of the Kim family. Correctness: correct Question: Which era did Michael Oakeshott belong to? Ground truth: ["20th-century philosophy"] Prediction: 20th century." Correctness: correct Question: Edward Tise (known for Full Metal Jacket (1987)) is in what department? Ground truth: ["sound department"] Prediction: 2nd Infantry Division, United States Army Correctness: incorrect Question: What wine region is Finger Lakes AVA a part of? Ground truth: ["New York wine"] Prediction: Finger Lakes AVA Correctness: incorrect Question: question
Ground truth: gold answer
Prediction: generated answer
Correctness: